[Brahms-l] Shift report 20050430 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Apr 30 2005 - 08:00:03 EDT
Shifter leader was FR
Shifters were 



 - Beam on since 2:50 am.
                                                                                
 - Physics data with FS @ 8deg-608B and MRS @ 45deg-1050A
   5 runs 14343-14347, each ~1 hour, with a total of 325k trig2 events.
   This brings the total trig2 for this setting to 355k.
                                                                                
 - No detector problem


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Apr 30 08:00:27 2005

This archive was generated by hypermail 2.1.8 : Sat Apr 30 2005 - 08:00:34 EDT